A probability denisty function for the random variable X is given by pi=ki, where k is a constant. What value must k be if X takes on integer values between 0 and 16, inclusive? Number (Note: three decimal places is sufficient)
Answer #1

here for this to be valid: must be equal to 1

=k*(0+1+2+3+4.....+16)=k*136 =1

k=1/136 =0.007